# Gas-phase metallicity gradients in galaxies at Key Words.: Galaxies: high-redshift ā Galaxies: abundances ā Galaxies: ISM ā Galaxies: evolution ā Techniques: imaging spectroscopy ā Techniques: high angular resolution1 Introduction

### v1

@@
The study of gas-phase metallicity and its spatial distribution at high redshift is crucial to understand the processes that shaped the growth and evolution of galaxies in the early Universe. Here we study the spatially resolved metallicity in three systems at 6ā8, namely A2744-YD4, BDF-3299, and COSMOS24108, with JWST NIRSpec IFU low-resolution ( 100) spectroscopic observations. These are among the highest- sources in which metallicity gradients have been probed so far. Each of these systems hosts several spatial components in the process of merging within a few kiloparsecs, identified from the rest-frame UV and optical stellar continuum and ionised gas emission line maps. The sources have heterogeneous properties, with stellar masses log() 7.6ā9.3, star formation rates (SFRs) 1ā15 Ā yr, and gas-phase metallicities 12+log(O/H) 7.7ā8.3, which exhibit a large scatter within each system. Their properties are generally consistent with those of the highest-redshift samples to date ( 3ā10), though the sources in A2744-YD4 and COSMOS24108 are at the high end of the mass-metallicity relation (MZR) defined by the 3ā10 sources. Moreover, the targets in this work follow the predicted slope of the MZR at 6ā8 from most cosmological simulations. The gas-phase metallicity gradients are consistent with being flat in the main sources of each system. Flat metallicity gradients are thought to arise from gas mixing processes on galaxy scales, such as mergers or galactic outflows and SN winds driven by intense stellar feedback, which wash out any gradient formed in the galaxy. The existence of flat gradients at 6ā8 sets also important constraints on cosmological simulations, whose predictions on the cosmic evolution of metallicity gradients often differ significantly, especially at high redshift.
@@
Negative (radially decreasing) gradients are usually interpreted as resulting from an inside-out galaxy formation scenario, in which stars start forming earlier in the inner parts of galaxies and thus have more time to chemically enrich the inner regions than the outer ones (e.g. Samland etĀ al., 1997; Prantzos & Boissier, 2000; DavĆ© etĀ al., 2011; Pilkington etĀ al., 2012; Gibson etĀ al., 2013; Hemler etĀ al., 2021; Tissera etĀ al., 2022). Flattened gradients could arise from radial mixing of gas and redistribution of metals on galaxy scales, induced by SN winds within the galaxy or galactic outflows of metal-enriched material expelled from the galaxy and re-accreted in the outer regions (so-called āgalactic fountainsā), driven by intense stellar feedback (e.g. Gibson etĀ al., 2013; Ma etĀ al., 2017), as well as by galaxy merging and interactions (e.g. Rupke etĀ al., 2010b, a; Rich etĀ al., 2012; Torres-Flores etĀ al., 2014). Finally, positive (radially increasing; also called āinvertedā) gradients may be produced by accretion of external pristine gas towards the central regions of the galaxy (e.g. Ceverino etĀ al., 2016); this could also contribute to flatten the metallicity gradient, if not invert it. Particularly strong metal-loaded galactic outflows could also contribute to produce a positive gradient, by moving the metal-rich gas from the central starburst regions to the outskirts (e.g. Tissera etĀ al., 2022). A schematic cartoon that summarises the above framework is shown in Fig. 1.
@@
Cosmological simulations of galaxy evolution make very different predictions on metallicity gradients at high redshift (up to 3), which can strongly vary depending on prescriptions on star formation and stellar feedback models and on their relative contribution to gas mixing as compared to mergers (e.g. Gibson etĀ al., 2013; Ma etĀ al., 2017; Hemler etĀ al., 2021; Tissera etĀ al., 2022). For example, according to the FIRE simulations (Hopkins etĀ al., 2014), mergers and rapid variations of metallicity gradients in high- galaxies, induced by starburst episodes which drive strong outflows and disrupt the gas disc, would flatten any negative gradients previously developed, especially in low-mass galaxies where feedback mechanisms are more efficient (Ma etĀ al., 2017). Instead, the TNG50 simulations (Pillepich etĀ al., 2019; Nelson etĀ al., 2019) which feature less bursty feedback, predict negative gradients at all redshifts, steeper (more negative) with increasing (Hemler etĀ al., 2021). The metallicity gradients resulting from the MUGS (ānormalā feedback; Stinson etĀ al., 2010) and MaGICC (āenhancedā feedback; Brook etĀ al., 2012) simulation suites are compared in Gibson etĀ al. (2013). The āenhancedā feedback (including pre-SN early stellar feedback from massive stars) re-distributes energy and re-cycled ISM material over large scales, through re-accretion in the outer parts of the galaxy of gas expelled via outflows. In this case, relatively flat and temporally invariant abundance gradients are predicted, in contrast to the steeper negative gradients at increasing emerging from the normal feedback scenario. Finally, the EAGLE simulations (Schaye etĀ al., 2015) find that the median gradient is zero (flat) at all redshifts, but with the scatter around the median increasing with due to individual galaxies transitioning between steep (either negative or positive) and flatter gradients (Tissera etĀ al., 2022). The authors report that this behaviour results from the higher frequency of major mergers at higher leading to episodes of enhanced accretion of low-metallicity gas which trigger intense star formation and ejection of metal-enriched gas.
@@
Observations at the cosmic noon epoch (i.e., 1ā4) have found heterogeneous results, with both negative, positive, and flat metallicity gradients (e.g. Cresci etĀ al., 2010; Yuan etĀ al., 2011; Queyrel etĀ al., 2012; Swinbank etĀ al., 2012; Stott etĀ al., 2014; Troncoso etĀ al., 2014; Jones etĀ al., 2013, 2015; Leethochawalit etĀ al., 2016; Wuyts etĀ al., 2016; Wang etĀ al., 2017, 2022; Carton etĀ al., 2018; Fƶrster Schreiber etĀ al., 2018; Curti etĀ al., 2020a). Nevertheless, the majority of high- measurements is consistent with little or no cosmic evolution of metallicity gradients, which are found to be flat or only moderately negative or positive ( dexĀ kpc; see e.g. the compilation in Curti etĀ al. 2020a).
@@
The James Webb Space Telescope (JWST) has recently opened up the possibility of measuring metallicity gradients in the first 1ā2 Gyr of the Universe by using rest-frame optical emission lines (Wang etĀ al., 2022; RodrĆguez Del Pino etĀ al., 2023; Arribas etĀ al., 2023) which are extensively used and calibrated at lower redshifts (e.g. Curti etĀ al., 2017, 2020b).
In this work, we study the spatially resolved gas-phase metallicity and investigate the shape of metallicity gradients at high redshift ( 6), by tracing the warm (Ā Ā Ā K) ionised gas emission. Specifically, we observed three high- systems with the Near-InfraRed Spectrograph (NIRSpec) on board JWST in its Integral Field Unit (IFU) mode with the low-spectral resolution PRISM/CLEAR (āā100) disperser-filter combination. These are A2744-YD4, part of the proto-cluster A2744-z7p9OD (Ā Ā 7.88), BDF-3299 (Ā Ā 7.11), and COSMOS24108 (Ā Ā 6.36). Basic information on the targets is given in Table 1. We trace gas-phase metallicity by making use of strong-line calibrators relying on rest-frame optical and near-UV emission lines. We also present the basic properties of each source, specifically stellar mass, obtained from SED fitting, and SFR, from H or H. The star formation history of these targets is studied in detail in a separate paper (Kohandel et al., in prep.).
Throughout this work, the reported wavelengths are in vacuum and we adopt a flat CDM cosmology with 67.7Ā kmĀ sĀ Mpc, 0.31, and 0.69 (Planck Collaboration etĀ al., 2020).
@@ 2 Description of observations, data reduction, and analysis
The observations employed in this work were carried out between October and December 2022 as part of the JWST GO program ID 1893 (PI: S. Carniani; Carniani etĀ al. 2021). The three sources were observed for 15.102 ks (4 h) on-source each (22.059 ks with overheads; 6 h) with 1.678 ks of background exposure time each (2.830 ks with overheads). The observations were performed with the PRISM/CLEAR disperser-filter combination and an eight-point dither āmediumā cycling pattern (step 0.5). The JWST NIRSpec 3 IFU PRISM observations simultaneously span the spectral range 0.6ā5.2Ā m with a spectral resolution ranging between 30ā330.
Target name | RA [hh:mm:ss.ss] | Dec [dd:mm:ss.ss] | a | Scale [kpc arcsec] |
A2744-YD4 | +00:14:24.90 | ā30:22:56.10 | 7.879 | 3.512b |
@@ 2.1 Data reduction
We retrieved the raw data from the MAST archive and we ran the three stages of the pipeline using version 1.11 with CRDS (calibration reference data system) context ājwst_1094.pmapā. First, at stage 1 ācalwebb_detector1ā, the pipeline applied the detector-level corrections (e.g. check for saturation, dark exposure subtraction, flagging of bad pixels and cosmic-ray persistences) and performed ramp fitting for individual exposures. We then calibrated the count rate images by executing stage 2 ācalwebb_spec2ā of the pipeline, which corrects for flat field and performs the wavelength calibration. The background was subtracted from each exposure during this stage by using the observations of the dedicated background for each target. We processed the background targets up to stage 2 of the pipeline, then we applied the background step for the science target. In the background step, the pipeline subtract the background exposure from each target exposure in the detector space. Finally, each calibrated exposure was combined in stage 3 ācalwebb_spec3ā by using ādrizzleā weighting and a spaxel size of 0.05ā³Ā to obtain the final cubes. During stage 3, we applied the outlier rejection step built in the pipeline and then a sigma clipping to remove any residual outliers in the final cube.
@@ 2.2 Data analysis
R | [Oāii] 3727,30/H |
R | [Oāiii] 5008/H |
OO | [Oāiii] 5008/[Oāii] 3727,30 |
0.47 R + 0.88 R | |
NeO | [Neāiii] 3870/[Oāii] 3727,30 |
RS | [Oāiii] 5008/HĀ + [Sāii] 6718,33/H |
OS | [Oāiii] 5008/HĀ / [Sāii] 6718,33/H |
In this section we describe the analysis of the NIRSpec R100 IFU data. In brief, we obtained the emission line fluxes from emission-line modelling of the spectra, which were used to infer the gas-phase metallicity. This was done on both an integrated basis (to get the integrated metallicity of each target) and a spatially resolved one, in this case both spaxel-by-spaxel (to produce maps) and in concentric radial annuli (to obtain radial gradients). From the integrated spectra, we also obtained the stellar mass () of each source from SED fitting and their star formation rate (SFR) from the emission-line modelling. We provide more details in the following.
@@ 2.2.1 Emission line fitting
The main goal of this work is to measure radial gas-phase metallicity gradients. To do so, we infer the oxygen abundance relative to hydrogen (12+log(O/H)), a proxy of gas-phase metallicity, by making use of the optical and near-UV strong-line diagnostic ratios reported in Table 2. We adopt the new diagnostic ratio = 0.47 R + 0.88 R, first introduced in Laseter etĀ al. (2024), in place of the more traditional R = ([Oāii] 3727,30 + [Oāiii] 4960,5008)/H, since the former is more suited for high- galaxies, while the latter is a projection mostly driven by low-metallicity local analogues. We adopt the best-fit polynomial calibrations from Curti etĀ al. (2017, 2020b), slightly revisited in Curti etĀ al. (2023a, b) to better probe the low-O/H regime, to infer the gas metallicity from the combination of the above ratios. We tested other gas-phase metallicity calibrations (e.g. Nakajima etĀ al., 2022; Sanders etĀ al., 2024). These gave 12+log(O/H) values similar to the Curti et al. ones within 0.1 dex, consistent with the uncertainties, and virtually no difference in the metallicity radial gradients.
@@
The FWHM of the point spread function (PSF) of JWST depends on wavelength. Specifically, the PSF FWHM of NIRSpec IFU ranges from around 0.09 at 1 m up to around 0.16 at 4.5 m, and the variation with wavelength is stronger in the direction perpendicular to the IFU slices than along them (DāEugenio etĀ al., 2023). Since the aim of this work is to obtain spatially resolved metallicities from emission line ratios sometimes far in wavelength (see Table 2), we applied a wavelength-dependent spatial smoothing to the data cube prior to the line fitting with the aim of achieving the same spatial resolution at all wavelengths. We adopted the PSF FWHM curves from DāEugenio etĀ al. (2023), empirically calibrated by matching NIRSpec IFU observations with NIRCam ones of the same target reported in their Eqs. 3 and 4, for the along- and across-slice cases, respectively. The smoothing was done by convolving the cube with a wavelength-dependent 2D Gaussian kernel in order to obtain a smoothed cube having the same spatial resolution at each wavelength, specifically the PSF FWHM reported in DāEugenio etĀ al. (2023) at the wavelength of the highest-wavelength line of interest for the along-slice case (which has the largest FWHM among the two cases). This corresponds to FWHM 0.16 (at [Oāiii]) for A2744-YD4, 0.14 (at [Oāiii]) for BDF-3299, and 0.18 (at H-[Sāii]) for COSMOS24108. Specifically, the 2D Gaussian kernel to be used for the convolution, at a given wavelength , was defined as follows: FWHM() = FWHM() ā FWHM(), where FWHM is the model FWHM from DāEugenio etĀ al. (2023) and ādirā indicates the direction, either along or across the slices.
@@
In general, we employ the original unsmoothed data cube for the flux maps of continuum and emission lines and for integrated measurements from circular apertures, while we adopt the smoothed data cube for the maps and radial profiles involving line ratios and metallicity, which would otherwise be affected by the wavelength-dependence of the PSF. In case of A2744-YD4, the highest- target of our sample, we adopt the smoothed data cube for every map, to obtain visually clearer maps as compared to the more noisy ones from the unsmoothed cube.
@@
The data analysis consisted in modelling the rest-frame optical and near-UV emission lines available in the observed spectral range. The emission lines are spectrally unresolved in the PRISM spectra (the spectral resolution is 1250 kmĀ s at 3 m and 450 kmĀ s at 5 m), therefore a single Gaussian function per line was used. We fitted the [Oāiii]-H, the [Oāii]-[Neāiii]-H, and the H-[Nāii]-[Sāii] line complexes separately. The [Nāii] 6550,85 was included to allow for a better modelling of the H line profile, which was otherwise showing a marked residual in its redward wing. This was done only when fitting integrated spectra from circular apertures or concentric annuli for radial profiles, not in the lower-S/N case of spaxel-by-spaxel fitting, for which no residual wing indicative of [Nāii] was detected above the noise. For each line complex, the velocity was tied to be the same for all the lines, while we allowed the line width to vary to match the PRISM spectral resolution, given that all lines are spectrally unresolved as mentioned. We fixed the flux ratios [Oāiii] 5008/4960 and [Nāii] 6585/6550 to their theoretical value of 3 (Storey & Zeippen, 2000). We included an underlying first-order polynomial to model the continuum. We accounted for the Balmer break of the continuum at around 3645 Ć Ā rest-frame, when needed, by employing three (first-order) polynomials, one modelling the jump and the other two on each side of it.
@@
The line fluxes used to infer the metallicity or the SFR were first corrected for dust extinction, when possible (see below), using the H/H ratio (for A2744-YD4 and BDF-3299) or H/H (for COSMOS24108), only when both lines had S/N3. We adopted a Calzetti etĀ al. (2000) reddening curve, with an = 4.05, suitable for high- low-metallicity (12+log(O/H) 8.5) star-forming galaxies at the wavelengths of interest here of 3000Ā Ć Ā (see e.g. Shivaei etĀ al. 2020). We assumed the theoretical extinction-free H/H and H/H ratios of 0.466 and 2.87, respectively, valid for case-B recombination and an electron temperature of 10Ā K (Osterbrock & Ferland, 2006), typical of the warm ionised gas emitting rest-frame optical lines.
@@
The emission line modelling was then performed both on a spaxel-by-spaxel basis (for both the unsmoothed and the smoothed data cubes) and in concentric radial annuli centred on each source (for the smoothed data cube). The latter was done to increase the S/N on the emission lines and get more robust estimates of metallicity with the goal of obtaining metallicity gradients. For this annular line modelling, we extracted integrated spectra at each radius by collapsing the spaxels within concentric circular annuli. The annuli, having radial width of 1 spaxel (0.05) each, were centred on each of the main spatial components detected in each targeted system, with a variable maximum aperture radius depending on the component extension (as described in the next sections separately for each target).

### v2

@@
The study of gas-phase metallicity and its spatial distribution at high redshift is crucial to understand the processes that shaped the growth and evolution of galaxies in the early Universe. Here we study the spatially resolved metallicity in three systems at 6ā8, namely A2744-YD4, BDF-3299, and COSMOS24108, with JWST NIRSpec IFU low-resolution ( 100) spectroscopic observations. These are among the highest- sources in which metallicity gradients have been probed so far. Each of these systems hosts several spatial components in the process of merging within a few kiloparsecs, identified from the rest-frame UV and optical stellar continuum and ionised gas emission line maps. The sources have heterogeneous properties, with stellar masses log() 7.6ā9.3, star formation rates (SFRs) 1ā15 Ā yr-1, and gas-phase metallicities 12+log(O/H) 7.7ā8.3, which exhibit a large scatter within each system. Their properties are generally consistent with those of the highest-redshift samples to date ( 3ā10), though the sources in A2744-YD4 and COSMOS24108 are at the high end of the mass-metallicity relation (MZR) defined by the 3ā10 sources. Moreover, the targets in this work follow the predicted slope of the MZR at 6ā8 from most cosmological simulations. The gas-phase metallicity gradients are consistent with being flat in the main sources of each system. Flat metallicity gradients are thought to arise from gas mixing processes on galaxy scales, such as mergers or galactic outflows and supernova winds driven by intense stellar feedback, which wash out any gradient formed in the galaxy. The existence of flat gradients at 6ā8 sets also important constraints on future cosmological simulations and chemical evolution models, whose predictions on the cosmic evolution of metallicity gradients often differ significantly, especially at high redshift, but are mostly limited to 3 so far.
@@
Negative (radially decreasing) gradients are usually interpreted as resulting from an inside-out galaxy formation scenario, in which stars start forming earlier in the inner parts of galaxies and thus have more time to chemically enrich the inner regions than the outer ones (e.g. Samland et al., 1997; Prantzos & Boissier, 2000; Davé et al., 2011; Pilkington et al., 2012; Gibson et al., 2013; Hemler et al., 2021; Tissera et al., 2022). Flattened gradients could arise from radial mixing of gas and redistribution of metals on galaxy scales, induced by supernova (SN) winds within the galaxy or galactic outflows of metal-enriched material expelled from the galaxy and re-accreted in the outer regions (so-called galactic fountains), driven by intense stellar feedback (e.g. Gibson et al., 2013; Ma et al., 2017), as well as by galaxy merging and interactions (e.g. Rupke et al., 2010b, a; Rich et al., 2012; Torres-Flores et al., 2014). Finally, positive (radially increasing; also called inverted) gradients may be produced by accretion of external pristine gas towards the central regions of the galaxy (e.g. Ceverino et al., 2016); this could also contribute to flatten the metallicity gradient, if not invert it. Particularly strong metal-loaded galactic outflows could also contribute to produce a positive gradient, by moving the metal-rich gas from the central starburst regions to the outskirts (e.g. Tissera et al., 2022). A schematic cartoon that summarises the above framework is shown in Fig. 1.
@@
Cosmological simulations of galaxy evolution make very different predictions on metallicity gradients at high redshift (up to 3), which can strongly vary depending on prescriptions on star formation and stellar feedback models and on their relative contribution to gas mixing as compared to mergers (e.g. Gibson etĀ al., 2013; Ma etĀ al., 2017; Hemler etĀ al., 2021; Tissera etĀ al., 2022). For example, according to the FIRE simulations (Hopkins etĀ al., 2014), mergers and rapid variations of metallicity gradients in high- galaxies, induced by starburst episodes which drive strong outflows and disrupt the gas disc, would flatten any negative gradients previously developed, especially in low-mass galaxies where feedback mechanisms are more efficient (Ma etĀ al., 2017). Instead, the TNG50 simulations (Pillepich etĀ al., 2019; Nelson etĀ al., 2019) which feature less bursty feedback, predict negative gradients at all redshifts, steeper (more negative) with increasing (Hemler etĀ al., 2021). The metallicity gradients resulting from the MUGS (ānormalā feedback; Stinson etĀ al., 2010) and MaGICC (āenhancedā feedback; Brook etĀ al., 2012) simulation suites are compared in Gibson etĀ al. (2013). The enhanced feedback (including pre-SN early stellar feedback from massive stars) re-distributes energy and re-cycled ISM material over large scales, through re-accretion in the outer parts of the galaxy of gas expelled via outflows. In this case, relatively flat and temporally invariant abundance gradients are predicted, in contrast to the steeper negative gradients at increasing emerging from the normal feedback scenario. Finally, the EAGLE simulations (Schaye etĀ al., 2015) find that the median gradient is zero (flat) at all redshifts, but with the scatter around the median increasing with due to individual galaxies transitioning between steep (either negative or positive) and flatter gradients (Tissera etĀ al., 2022). The authors report that this behaviour results from the higher frequency of major mergers at higher leading to episodes of enhanced accretion of low-metallicity gas which trigger intense star formation and ejection of metal-enriched gas.
Non-cosmological chemical evolution models have also been employed to study the evolution of metallicity gradients with cosmic time, mostly for the Milky Way. Mott etĀ al. (2013), assuming an inside-out formation of the disc, a constant star formation efficiency (SFE) along the disc, and the presence of radial flows with varying speed, predict an inversion of the gradients (from negative to positive) at 1, due to a strong infall of primordial gas in the innermost disc regions at early times; only a variable SFE does not lead to an inversion of the gradients at high in their models. Kubryk etĀ al. (2015) study the role of the radial motions of gas and stars on the evolution of abundance profiles in the Milky Way disc, finding steep abundance profiles at high which flatten with time, as a result of the inside-out formation of the disc. MollĆ” etĀ al. (2019) investigate the role of the growth of the stellar disc, the effect of infall rate and star formation prescriptions, as well as the pre-enrichment of the infall gas, and find a smooth evolution of the gradients with a slight flattening from = 4 to 1. Finally, the first principles-based model of Sharda etĀ al. (2021) finds that the gradient in Milky Way-like galaxies has steepened over time and also predicts the evolution of metallicity gradients with redshift in galaxy samples matched in both stellar masses and abundances, finding that disc galaxies transition from the advection- to the accretion-dominated regime from high to low ; in general little evolution of the gradients is predicted for 1.
@@
Observations at the cosmic noon epoch (i.e., 1ā4) have found heterogeneous results, with both negative, positive, and flat metallicity gradients (e.g. Cresci etĀ al., 2010; Yuan etĀ al., 2011; Queyrel etĀ al., 2012; Swinbank etĀ al., 2012; Stott etĀ al., 2014; Troncoso etĀ al., 2014; Jones etĀ al., 2013, 2015; Leethochawalit etĀ al., 2016; Wuyts etĀ al., 2016; Wang etĀ al., 2017, 2022; Carton etĀ al., 2018; Fƶrster Schreiber etĀ al., 2018; Curti etĀ al., 2020a). Nevertheless, the majority of high- measurements is consistent with little or no cosmic evolution of metallicity gradients, which are found to be flat or only moderately negative or positive ( dexĀ kpc-1; see e.g. the compilation in Curti etĀ al. 2020a).
@@
The James Webb Space Telescope (JWST) has recently opened up the possibility of measuring metallicity gradients in the first 1ā2 Gyr of the Universe by using rest-frame optical emission lines (Wang etĀ al., 2022; RodrĆguez Del Pino etĀ al., 2023; Arribas etĀ al., 2023) which are extensively used and calibrated at lower redshifts (e.g. Curti etĀ al., 2017, 2020b). In this work, we study the spatially resolved gas-phase metallicity and investigate the shape of metallicity gradients at high redshift ( 6), by tracing the warm (Ā Ā Ā K) ionised gas emission. Specifically, we observed three high- systems with the Near-InfraRed Spectrograph (NIRSpec) on board JWST in its Integral Field Unit (IFU) mode with the low-spectral resolution PRISM/CLEAR (āā100) disperser-filter combination. These are A2744-YD4, part of the proto-cluster A2744-z7p9OD (Ā Ā 7.88), BDF-3299 (Ā Ā 7.11), and COSMOS24108 (Ā Ā 6.36). Basic information on the targets is given in Table 1. We trace gas-phase metallicity by making use of strong-line calibrators relying on rest-frame optical and near-UV emission lines. We also present the basic properties of each source, specifically stellar mass, obtained from spectral energy distribution (SED) fitting, and star formation rate (SFR), from H or H. The star formation history of these targets is studied in detail in a separate paper (Kohandel et al., in prep.).
Throughout this work, the reported wavelengths are in vacuum and we adopt a flat CDM cosmology with 67.7Ā kmĀ s-1Ā Mpc-1, 0.31, and 0.69 (Planck Collaboration etĀ al., 2020).
@@ 2 Description of observations, data reduction, and analysis
The observations employed in this work were carried out between October and December 2022 as part of the JWST GO program ID 1893 (PI: S. Carniani; Carniani etĀ al. 2021). The three sources were observed for 15.102 ks (4 h) on-source each (22.059 ks with overheads; 6 h) with 1.678 ks of background exposure time each (2.830 ks with overheads). The observations were performed with the PRISM/CLEAR disperser-filter combination and an eight-point dither āmediumā cycling pattern (step 0.5ā²ā²). The JWST NIRSpec 3 IFU PRISM observations simultaneously span the spectral range 0.6ā5.2Ā m with a spectral resolution ranging between 30ā330.
[…]
@@ 2.1 Data reduction
We retrieved the raw data from the MAST archive and we ran the three stages of the pipeline using version 1.11 with CRDS (calibration reference data system) context ājwst_1094.pmapā. First, at stage 1 ācalwebb_detector1ā, the pipeline applied the detector-level corrections (e.g. check for saturation, dark exposure subtraction, flagging of bad pixels and cosmic-ray persistences) and performed ramp fitting for individual exposures. We then calibrated the count rate images by executing stage 2 ācalwebb_spec2ā of the pipeline, which corrects for flat field and performs the wavelength calibration. The background was subtracted from each exposure during this stage by using the observations of the dedicated background for each target. We processed the background targets up to stage 2 of the pipeline, then we applied the background step for the science target. In the background step, the pipeline subtracts the background exposure from each target exposure in the detector space. Finally, each calibrated exposure was combined in stage 3 ācalwebb_spec3ā by using ādrizzleā weighting and a spaxel size of 0.05ā³Ā to obtain the final cubes. During stage 3, we applied the outlier rejection step built in the pipeline and then a sigma clipping to remove any residual outliers in the final cube.
@@ 2.2 Data analysis
| R2 | [Oāii] 3727,30/H |
| R3 | [Oāiii] 5008/H |
| O3O2 | [Oāiii] 5008/[Oāii] 3727,30 |
| 0.47 R2 + 0.88 R3 | |
| Ne3O2 | [Neāiii] 3870/[Oāii] 3727,30 |
| RS32 | [Oāiii] 5008/HĀ + [Sāii] 6718,33/H |
| O3S2 | [Oāiii] 5008/HĀ / [Sāii] 6718,33/H |
In this section, we describe the analysis of the NIRSpec R100 IFU data. In brief, we obtained the emission line fluxes from emission-line modelling of the spectra, which were used to infer the gas-phase metallicity. This was done on both an integrated basis (to get the integrated metallicity of each target) and a spatially resolved one, in this case both spaxel-by-spaxel (to produce maps) and in concentric radial annuli (to obtain radial gradients). From the integrated spectra, we also obtained the stellar mass () of each source from SED fitting and their SFR from the emission-line modelling. We provide more details in the following.
@@ 2.2.1 Emission line fitting
The main goal of this work is to measure radial gas-phase metallicity gradients. To do so, we infer the oxygen abundance relative to hydrogen (12+log(O/H)), a proxy of gas-phase metallicity, by making use of the optical and near-UV strong-line diagnostic ratios reported in Table 2. We adopt the new diagnostic ratio = 0.47 R2 + 0.88 R3, first introduced in Laseter etĀ al. (2024), in place of the more traditional R23 = ([Oāii] 3727,30 + [Oāiii] 4960,5008)/H, since the former is more suited for high- galaxies, while the latter is a projection mostly driven by low-metallicity local analogues. We adopt the best-fit polynomial calibrations from Curti etĀ al. (2017, 2020b), slightly revisited in Curti etĀ al. (2023a, b) to better probe the low-O/H regime, to infer the gas metallicity from the combination of the above ratios. We tested other gas-phase metallicity calibrations (e.g. Nakajima etĀ al., 2022; Sanders etĀ al., 2024). These gave 12+log(O/H) values similar to the Curti et al. ones within 0.1 dex, consistent with the uncertainties, and virtually no difference in the metallicity radial gradients.
@@
The full width at half maximum (FWHM) of the point spread function (PSF) of JWST depends on wavelength. Specifically, the PSF FWHM of NIRSpec IFU ranges from around 0.09ā²ā² at 1 m up to around 0.16ā²ā² at 4.5 m, and the variation with wavelength is stronger in the direction perpendicular to the IFU slices than along them (DāEugenio etĀ al., 2023). Since the aim of this work is to obtain spatially resolved metallicities from emission line ratios sometimes far in wavelength (see Table 2), we applied a wavelength-dependent spatial smoothing to the data cube prior to the line fitting with the aim of achieving the same spatial resolution at all wavelengths. We adopted the PSF FWHM curves from DāEugenio etĀ al. (2023), empirically calibrated by matching NIRSpec IFU observations with NIRCam ones of the same target reported in their Eqs. 3 and 4, for the along- and across-slice cases, respectively. The smoothing was done by convolving the cube with a wavelength-dependent 2D Gaussian kernel in order to obtain a smoothed cube having the same spatial resolution at each wavelength, specifically the PSF FWHM reported in DāEugenio etĀ al. (2023) at the wavelength of the highest-wavelength line of interest for the along-slice case (which has the largest FWHM among the two cases). This corresponds to FWHMgoal 0.16ā²ā² (at [Oāiii]) for A2744-YD4, 0.14ā²ā² (at [Oāiii]) for BDF-3299, and 0.18ā²ā² (at H-[Sāii]) for COSMOS24108. Specifically, the 2D Gaussian kernel to be used for the convolution, at a given wavelength , was defined as follows: FWHM() = FWHM() ā FWHM(), where FWHMmod;dir is the model FWHM from DāEugenio etĀ al. (2023) and ādirā indicates the direction, either along or across the slices.
@@
In general, we employed the original unsmoothed data cube for the flux maps of continuum and emission lines and for integrated measurements from circular apertures, while we adopted the smoothed data cube for the maps and radial profiles involving line ratios and metallicity, which would have otherwise been affected by the wavelength-dependence of the PSF. In case of A2744-YD4, the highest- target of our sample, we adopted the smoothed data cube for every map, to obtain visually clearer maps as compared to the more noisy ones from the unsmoothed cube.
@@
The data analysis consisted in modelling the rest-frame optical and near-UV emission lines available in the observed spectral range. The emission lines are spectrally unresolved in the PRISM spectra (the spectral resolution is 1250 kmĀ s-1 at 3 m and 450 kmĀ s-1 at 5 m), therefore a single Gaussian function per line was used. We fitted the [Oāiii]-H, the [Oāii]-[Neāiii]-H, and the H-[Nāii]-[Sāii] line complexes separately. The [Nāii] 6550,85 was included to allow for a better modelling of the H line profile, which was otherwise showing a marked residual in its redward wing. This was done only when fitting integrated spectra from circular apertures or concentric annuli for radial profiles, not in the lower-S/N case of spaxel-by-spaxel fitting, for which no residual wing indicative of [Nāii] was detected above the noise. For each line complex, the velocity was tied to be the same for all the lines, while we allowed the line width to vary to match the PRISM spectral resolution, given that all lines are spectrally unresolved as mentioned. We fixed the flux ratios [Oāiii] 5008/4960 and [Nāii] 6585/6550 to their theoretical value of 3 (Storey & Zeippen, 2000). We included an underlying first-order polynomial to model the continuum. We accounted for the Balmer break of the continuum at around 3645 Ć Ā rest-frame, when needed, by employing three (first-order) polynomials, one modelling the jump and the other two on each side of it.
@@
The line fluxes used to infer the metallicity or the SFR were first corrected for dust extinction, when possible (see below), using the H/H ratio (for A2744-YD4 and BDF-3299) or H/H (for COSMOS24108), only when both lines had S/N3. We adopted a Calzetti etĀ al. (2000) reddening curve, with an = 4.05, suitable for high- low-metallicity (12+log(O/H) 8.5) star-forming galaxies at the wavelengths of interest here of 3000Ā Ć Ā (see e.g. Shivaei etĀ al. 2020). We assumed the theoretical extinction-free H/H and H/H ratios of 0.466 and 2.87, respectively, valid for case-B recombination and an electron temperature of 104Ā K (Osterbrock & Ferland, 2006), typical of the warm ionised gas emitting rest-frame optical lines.
@@
The emission line modelling was then performed both on a spaxel-by-spaxel basis (for both the unsmoothed and the smoothed data cubes) and in concentric radial annuli centred on each source (for the smoothed data cube). The latter was done to increase the S/N on the emission lines and get more robust estimates of metallicity with the goal of obtaining metallicity gradients. For this annular line modelling, we extracted integrated spectra at each radius by collapsing the spaxels within concentric circular annuli. The annuli, having radial width of 1 spaxel (0.05ā²ā²) each, were centred on each of the main spatial components detected in each targeted system, with a variable maximum aperture radius depending on the component extension (as described in the next sections separately for each target).